Find the phase difference between two points having phase difference 2pi radians
 
 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Assuming the question is-

Find the path difference between two points having phase difference 2pi radian.

The path difference between points is\dfrac{\lambda}{2\pi}\times 2\pi=\lambda